You wish to produce a potential difference of 40 V along the length of a 1 m copper rod by moving it through the earth’s magnetic field. In a region where the field is 5.5×10−5T, what minimum speed would be needed?
Answer #1

We know that

Potential difference or emf = vBLsin(theta)

here v = velocity =?

B = Magnetic field = 5.5×10−5T

L = length of wire = 1 m

theta = 90 deg

voltage V = 40 V

hence velocity v = V/BL = 40/ 5.5×10−5 = 7,27,272.72 m/s...............Ans.